OpenStudy (anonymous):

-2ab/a^2-2ac a=2, b=-2, c=-3 (correct 1/6?)

OpenStudy (anonymous):

I come up with 8/48 reduced to 1/6...not sure tho

OpenStudy (angela210793):

i got 1/2

OpenStudy (anonymous):

what is the number in the deno u got to get two?

OpenStudy (akshay_budhkar):

check the denominator again

OpenStudy (angela210793):

i got 8/16=1/2

OpenStudy (anonymous):

still not seeing the 16...

OpenStudy (anonymous):

2^2-2(2)(-3) how is it 16??

OpenStudy (angela210793):

4-4*(-3)=4+12=16

OpenStudy (akshay_budhkar):

get it?

OpenStudy (angela210793):

who me? I do

OpenStudy (akshay_budhkar):

lol not you!!!!!!!

OpenStudy (anonymous):

trying...ty!!!

OpenStudy (akshay_budhkar):

you are a MASTER

OpenStudy (angela210793):

who me? I do

OpenStudy (angela210793):

lool :D ok...so I'm right

OpenStudy (akshay_budhkar):

i have no right as a HERO to ask you angela :)

OpenStudy (angela210793):

Lool..:D :P

OpenStudy (anonymous):

i need to work left to right...;)

OpenStudy (akshay_budhkar):

you still working? whats that u not getting?

OpenStudy (anonymous):

I think just the 4-4(-3), I either want to cancel the first two out or multiply all together

OpenStudy (angela210793):

but u multiply 1st then add or subtract this is how it works

OpenStudy (anonymous):

But I see it once I typed it....I have to subtract that first...